Skip to content

VTRx OH129

On April 19-20th of 2021, the RSSI current was monitored for the VTRx in GBT0, GBT1, and GBT2, of OH129 in detector GE11-X-S-FIT-0002 (Bottom layer SCS027 extracted from P-17@P5). After 24h of monitoring, the RSSI currents were not stable, instead of that, VTRx0 and VTRx1 fell below 50 uA.

Then the three VTRx were replaced: VTRx on GBT0 → removed 209764 , Inserted 209778 VTRx on GBT1 → removed 215238 , Inserted 209779 VTRx on GBT0 → removed 209237 , Inserted 209780

Initially the RSSI currents observed for the new set of VTRx were very low (~60 uA instead of 300uA ), but after cleaning the fibers carefully, RSSI values became normal.

The new set of VTRx, showed stable RSSI currents in 24 hours of monitoring.

rssi_FIT0002S_20210419_snrssi_FIT0002S_2021.04.20_sn

Edited by Johny Jaramillo
To upload designs, you'll need to enable LFS and have an admin enable hashed storage. More information